### v1

# Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

Abstract
In Production Scheduling, the Flexible Job Shop Scheduling Problem (FJSSP) aims to optimize a sequence of operations and assign each to an eligible machine with varying processing times. To be even closer to real production environments, the integration of the workforce may be considered. Hence, each machine also requires a worker to be present to process an operation which additionally affects the processing times. The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
The FJSSP has been approached with different strategies considering various problem representations, including Mixed Integer Linear Programming (MILP), Constrained Programming (CP), and Simulation-based Optimization (SBO). In the latter area in particular, there exists a large number of specialized Evolutionary Algorithms (EA) like Particle Swarm Optimization (PSO) or Genetic Algorithms (GA). Yet, the solvers are often developed for single use cases only, and validated on a few selected test instances, let alone compared with results from state-of-the-art solvers using other problem representations. While suitable approaches do also exist, the design of the FJSSP-W instances is not standardized and the performance of the algorithms is hardly comparable.
This calls for a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances and supports targeted algorithm development. It will facilitate the comparison of algorithmic performance in the face of different problem characteristics. The present paper presents a collection of 402 commonly accepted FJSSP instances and proposes an approach to extend these with worker flexibility. The conversion of the FJSSP into an FJSSP-W has a significant impact on problem complexity. In addition, we present a detailed procedure for the evaluation of scheduling algorithms on these problem sets and provide suitable model representations for this purpose. This includes MILP formulations, CP models, and appropriate problem encodings for SBO solvers for both problem types. In addition, we provide complexity characteristics for all presented instances as well as baseline results of common commercial solvers to facilitate the validation of new algorithmic developments.
keywords:
Simulation-Based Optimization , Evolutionary Computation , Production Planning , Flexible Job Shop Scheduling , Worker Flexibility , Benchmarking
journal: Swarm and Evolutionary Computation
[…]
[inst1]organization=Josef Ressel Centre for Robust Decision Making
Research Centre Business Informatics
Vorarlberg University of Applied Sciences,addressline=
Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
{highlights}
An alternative modeling approach for flexible job shop scheduling problems (FJSSP) as a mixed integer linear program (MILP) suitable to represent an exhaustive set of commonly accepted benchmark problems is proposed.
The paper presents a refined algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a portfolio of benchmark instances for an otherwise rather fragmented problem branch. A corresponding extension of the derived MILP representation for the introduced FJSSP-W is also delivered.
We provide a suitable problem encoding for a convenient simulation-based optimization (SBO) via meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, to such problem classes and present an approach to effectively treat the worker flexibility constraints.
In addition, we develop adequate Constraint Programming (CP) representations for both problem classes (FJSSP and FJSSP-W) which facilitate the comparison with state-of-the-art solvers from this field of research.
Two problem classes in three representations (MILP, CP, and SBO) provide a comprehensive test environment for the comparison of solvers of different algorithmic paradigms. As a baseline, we compare the results of commercial state-of-the-art solvers (GUROBI, Hexaly, CPLEX (IBM), OR-Tools (Google)) and self-developed Genetic Algorithms (GA).
1 Introduction
Planning problems occur in a variety of application areas where the allocation and optimization of resources is critical to achieving specific goals. Consequently, planning problems are relevant in virtually all economic areas and the corresponding challenges are critical for ensuring sustainable economic growth, competitiveness, and efficient operations. In Supply Chain Management, for example, planning is essential for coordinating the flow of goods. In the Energy Sector, planning problems deal with the generation, distribution, and consumption of energy. Other application fields include Healthcare, Finance, Agriculture, and many more. Another important area is Manufacturing, i.e. the branch of industry that involves the production of finished products from raw materials. There, Production Scheduling (PS) deals with the allocation of resources such as machinery, labor and materials in order to achieve production targets efficiently. Such production goals include meeting customer requirements in terms of delivery times, minimizing idle time in certain production steps or responding to machine breakdowns. In this way, effective scheduling aims to promote sustainable development, drive innovation and ensure the resilience and competitiveness of manufacturing companies.
Production Scheduling problems involve determining the proper sequence of actions or decisions to attain (near) optimal performance on the shop floor. They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI) or Operations Research (OR). A broad overview of the PS landscape is provided in Fig. 1, in which the problem branches on which this work is based are indicated by the blue boxes.
Figure 1: Localization of the problem classes under consideration in the area of Production Scheduling. Notice that other problem types exist and the illustration does by no means claim to be comprehensive.
Among such tasks, the Job Shop Scheduling Problem (JSSP) is a type of scheduling problem where multiple jobs (or tasks) share a common set of resources but have different processing requirements and routes through the production facility. Each job may require different machines or processes, and the order in which the jobs are processed can affect the overall efficiency of the system. JSSPs may turn out very complex due to the variability in job processing times, resource constraints, and the need to optimize the sequence of operations. Unlike the traditional JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allows jobs to be processed on different machines. This provides more flexibility in resource allocation to those machines that possesses the required capabilities for a job. In this respect, the FJSSP represents an extension of the JSSP that considers a more dynamic and adaptable environment. The flexibility allows for more dynamic and adaptive scheduling, but it also introduces additional complexity to the scheduling problem, and optimization algorithms need to consider the various combinations of machine assignments to achieve an (near) optimal schedule. An additional layer of complexity to the scheduling process is added when worker-related dynamics are also considered within a FJSSP. Such problems are commonly referred to as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W). There are many restrictions that may arise when considering worker flexibility in FJSSP, e.g., consecutive working hours, breaks or different qualifications of workers.
Depending on the modeling approach used and the optimization technique envisaged, JSSP and FJSSP model representations come in different forms. Common problem formulations include mathematical graphs, such as Resource-Task-Network (RTN) representations, permutation-based representation, or mathematical programming (MP) formulations in the form of objective functions and constraints. The latter usually come in the form of Mixed-Integer Linear Programming (MILP) problems and can be treated by various optimization techniques. Formulations as simulation-based optimization (SBO) problems are also frequently found in the literature. Instead of an analytical representation of the problem, the objectives and dependencies of a production line are implemented in the form of a black-box environment that allows the evaluation of individual production plans. Constraint Programming (CP) formulations are another way of expressing production planning problems. CP enables the integration of constraints and their solution mechanisms into a programming language. The problem is formulated in declarative form, while the solution process is secondary from the users’ point of view. This task is performed by a specialized constraint solver.
For SBO representations common approaches comprise meta-heuristic algorithms that rely on nature-inspired working principles, such as Genetic Algorithm (GA), Ant Colony Optimization (ACO) or Particle Swarm Optimization (PSO) variants. Hybrid approaches that combine multiple optimization techniques to exploit complementary strengths are also regularly considered. An example would be the combination of a GA with a Local Search method. Focusing on the exploration of the neighborhood of a given solution, Local Search (LS) solvers are used to refine solutions obtained from other methods and to find locally optimal solutions. In case of a MILP or CP formulation of the FJSSP, specialized commercial mathematical programming solvers of e.g. GUROBI and CPLEX (IBM) or constraint programming solvers like OR-Tools(Google), Hexaly or CP of CPLEX respectively, are widely used to find (near) optimal solutions. These solvers obtain state-of-the-art solutions for many FJSSP instances. Still, the choice of solver depends on factors such as the size and complexity of the FJSSP instance, the desired solution quality, and the available computational resources. Researchers and practitioners often experiment with multiple approaches to find the most effective solution for a specific FJSSP instance.
While the performance of most solvers is assessed on individual real-world problem instances and compared to selected solver variants only, some scheduling approaches are validated on rather small sets of test problem instances. In contrast to the real-world problems considered, the majority of FJSSP test sets must be considered fairly academic. Nevertheless, recent literature offers access to a number of standard FJSSP benchmark instances being spread across individual publications and modeled in different ways. Further, the benchmark problems usually vary in their structure and complexity. A unified benchmarking framework which allows dedicated algorithm performance comparisons on FJSSPs does not exist to the best of our knowledge. Moreover, there are hardly any comparable benchmark instances for FJSSP-W. As the modeling of FJSSP-W requires a careful consideration of both machine and worker-related constraints. The type of model affects the selection of solvers able to handle the increased complexity introduced by flexible scheduling policies. This introduces novel objectives like not only optimizing production efficiency but also taking into account the necessities of the workforce.
This paper proposes a systematic test environment of FJSSP and FJSSP-W instances to benchmark state-of-the-art solvers. For this purpose, an alternative MILP modeling approach for FJSSPs suitable for incorporation of a comprehensive set of 402 commonly accepted benchmark problems is introduced. Our MILP formulation allows for an improvement of the schedules found by the commercial solver GUROBI on some of the benchmark instances. The paper presents an algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a way to systematize the fragmented area of FJSSP-W instances. It provides an opportunity to create a portfolio of benchmark instances for the rather fragmented area of FJSSP-W instances. Further, suitable SBO problem encodings convenient for the application of meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, as well as adequate CP representations are provided. As a result, a comprehensive test environment for the comparison of solvers from three different paradigms (MILP, CP, and SBO) on both problem classes (of 402 instances each) is established. Lastly, the paper describes a structured benchmark environment for the FJSSP(-W) which offers the possibility to choose among different problem characteristics, as well as the evaluation of a set of performance indicators (relative solution quality, wall clock time, etc.) and corresponding graphical visualizations. The test environment also provides characteristic attributes for each problem instance, such as the total number of operations, the flexibility of the machines or the problem-related duration variety, on the basis of which a targeted selection of sub-instances can be made. The functionalities are demonstrated by presenting baseline results of various state-of-the-art commercial solvers (GUROBI, Local Solver, CPLEX (IBM), OR-Tools (Google)) and a self-developed Genetic Algorithms (GA).
The remainder of this paper is structured as follows: In Section 2, a detailed description of the considered problem classes is provided. In addition, related work is discussed and available benchmarks as well as state-of-the-art solvers are presented. Section 3 deals with the formulation of the respective FJSSP representations. Next, the procedure for integration of worker flexibility is specified in Sec. 4. It realizes the transformation of an FJSSP instance into an FJSSP-W instance with just a few user-defined assumptions. Subsequently, the corresponding FJSSP-W formulations are specified. The detailed description of the benchmarking environment and its configuration options are the subject of Sec. 5. Afterwards, Section 6 presents a number of baseline results obtained by application of several state-of-the-art solvers. The contribution of this work and its impact on the field of research is discussed in summary in Sec. 7. Finally, Section 8 gives an outlook on future extensions of the presented benchmarking suite.
2 Problem Formulation & Related Work
The Flexible Job Shop Scheduling Problem (FJSSP) is the starting point for the following investigations. In this section, a formal definition of this problem is presented in Sec. 2.1 and current work on the FJSSP is outlined in Sec. 2.3. A large proportion of such research looks at very individual problems from real-world applications or only compares the developed solvers on individually selected benchmark instances. This complicates the performance comparison between the solvers significantly. In order to counteract this situation and to develop a structured benchmarking environment, Sec. 2.2 provides an overview of the available benchmark instances in the field of FJSSP and their characteristics.
2.1 Problem Formulation
The class of FJSSPs consists of NP-complete [1] combinatorial optimization problems that require to sequence a series of operations on a subset of available machines in such a way that a certain cost function $C$ is minimized. The machines are not necessarily homogeneous which allows for different processing times for each operation, depending on which machine it is assigned to.
In general formulation, the FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$. Each of these jobs $J_{i}\>(i=1,\dots,n)$ consists of a predetermined sequence of $n_{i}$ operations. To this end, the notation $O_{i,j}$ refers to the $j$th operation within job $i$. Each of these operations $O_{i,j}$ is performed by one of the admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$. If operation $O_{i,j}$ is assigned to a specific machine, this machine is denoted by $M_{i,j}$. The total number of operations in a FJSSP can thus be calculated as $\texttt{N}=\sum_{i=1}^{n}n_{i}$. Accordingly, the ratio of the total amount of operations and the number of jobs gives the average number of operations per job $\frac{\texttt{N}}{n}$. The constant processing time of operation $O_{i,j}$ on machine $M_{k}\in\mathcal{M}_{i,j}\>(k=1,\dots,m)$ is given as $T_{i,j,k}$. The minimum processing duration among all possible operation-machine combinations of a single problem instance is referred to as ${T}_{\min}$. Conversely, the maximum processing duration is ${T}_{\max}$. Their difference defines the so-called duration span $\Delta T={T}_{\max}-{T}_{\min}$. In common scenarios, the general representation of an FJSSP assumes that all orders are known and all machines are operational at time 00.
In the context of FJSSPs, a large number of objective functions are conceivable with respect to the optimization of which the appropriate schedule must be composed. The performance criteria include minimizing the makespan, minimizing tardiness or maximizing resource utilization, to name but a few. While the makespan, i.e. the total time required for the execution of all jobs of a problem instance, is one of the most important objective functions in literature and it can be used as an illustrative example in the remainder of this paper, the objective function is to be represented as generally as possible. This allows the subsequent inclusion of various optimization objectives both in the problem modeling and in the performance analysis within the proposed benchmarking environment. Since every maximization task can be formulated equivalently as a minimization problem, w.l.o.g. it can be assumed that a given cost function is to be minimized in full satisfaction of the production constraints of the FJSSP formulated above.
Consequently, the cost function represents a mapping of a certain schedule representation $\mathbf{y}\in\mathcal{Y}$ to a particular cost function value $C(\mathbf{y})\in\mathds{R}$, i.e. $C:\mathcal{Y}\to\mathds{R}$. Hence, one is interested in the search space vector $\mathbf{y}$ that realizes a minimal $C(\mathbf{y})$ value. Note, that concrete examples of a parameter vector encoding for $\mathbf{y}$ is depending on the modeling approach and will be provided below.
As it is common in the field of meta-heuristic optimization, the expression fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e. a feasible schedule, in the subsequent sections of this work. That is, solution candidates of higher fitness (or quality) correspond to solutions of lower cost function values. The quality of two admissible scheduling solutions can consequently be compared on the basis of their objective function values. Accordingly, lower cost function values $C(\mathbf{y})$ correspond to better production plans $\mathbf{y}\in\mathcal{Y}$. For a proper comparison of two specific schedules, an order relation $\prec_{c}$ is defined by
$\mathbf{y},\mathbf{z}\in\mathcal{Y}:\qquad\mathbf{y}\prec_{c}\mathbf{z}\>\ratio\Leftrightarrow\>C(\mathbf{y})<C(\mathbf{z}).$ (1)
Hence, the fitness of $\mathbf{y}$ would be rated higher than that of $\mathbf{z}$ in the case of Eq. (1). The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
2.2 Existing FJSSP Benchmarks
This section discusses the existing benchmark instances for the Flexible Job Shop Scheduling Problem. While there are already a number of test problem collections available for the FJSSP, these are often not systematically used for the evaluation of the developed solvers. Thus, the number of algorithms that are benchmarked on a large number of different FJSSP instances is relatively small [2]. A majority of solvers are rather evaluated on problem subgroups originating from distinct sources, which means that the results of such studies are always subject to a latent problem selection bias. A large number of papers also only consider individual real-world problem representations. The algorithms developed in this way are highly specialized and are rarely compared carefully with established solvers. This has two potential implications: on the one hand, the advantages of rather good solvers may remain unclear if only demonstrated in a very specific setting and secondly, incremental approaches may be overvalued if they are not compared with state-of-the-art methods.
Taking into account the structure and characteristics of the instances in such benchmark sources, one observes that the proposed test collections differ significantly from each other. Some provide a large number of very similarly structured instances, others provide fewer instances but greater differences in terms of problem complexity.
These design differences between the benchmark collections complicates the selection of reasonable test instances. For example, it is difficult to decide which instances of what collection should be chosen to evaluate an algorithm on a group of rather complex FJSSP instances. Especially on such instances, performance advantages of meta-heuristics over commercial MILP solvers could be expected. Of course, the answer could be to test the solvers on all available instances. Yet, it can often be useful to consider specific problem groups separately in order to focus on the individual strengths of a solver. In cases where the exact optimum is already known, i.e. for instances of relatively low complexity, the computing time required to achieve the optimum would be of greater interest. On the other hand, improvements of the solution quality are desirable as long as the exact optimum is unknown.
In this work, all $402$ distinct FJSSP instances of the ten test problem collections in Table 1 are considered. The table also indicates the number of instances (# instances) contained in each collections as well as the average number of operations per job (# avg. ops per job). The latter is within the range of $3$ and $19.5$, indicating the differences between the proposed benchmarking sources. Apart from the benchmark sources in Table 1, other collections with close relationships must be named. These are the problem collections in [3] and [4]. The design in [4] is limited to problem instances of the classical Job Shop Scheduling Problem (JSSP) and thus renounces the flexibility to run any of the jobs on any machine. Conversely, the instances in [3] allow parallel processing of some operations within a job. Taking these instances into account would therefore make it necessary to extend the models. Due to the special properties listed, both benchmark sets range outside the scope of this paper and are thus excluded. Conclusively, to the best of our knowledge, the benchmark sources referred to in Table 1 represent an extensive set of established test instances matching the definition of the FJSSP in Sec. 2.1.
Table 1: An overview of ten FJSSP benchmark problem collections from literature.
No. Source name # instances # avg. ops per job Reference
1 BehnkeGeiger 60 5,0 [5]
2 Brandimarte 15 8,6 [6]
3 ChambersBarnes 21 11,7 [7]
4 DPpaulli 18 19,5 [8]
5 Fattahi 20 3,0 [9]
6 Kacem 4 3,2 [10]
7 Hurink Edata 66 8,9 [11]
8 Hurink Sdata 66 8,9 [11]
9 Hurink Rdata 66 8,9 [11]
10 Hurink Vdata 66 8,9 [11]
The usual specification of the production environments is carried out for each individual FJSSP instance in the form of a text file, which shows all the necessary information about the dependencies in the flexible job shop. An example is provided in Fig. 2 to facilitate the comprehension of the specifications. There, the 11th instance of the FJSSP collection from [7] is used to describe the problem encoding.
The first row of such a file commonly contains two (optionally three) numbers: the first value determines the number of jobs $n$ and the second one specifies the number of machines $m$ in the FJSSP instance. An optional 3rd value may indicate the average number of machines per operation. Although the latter is not necessary for the implementation of the problem instance. In the example above, the number of jobs is $n=10$, there are $m=10$ machine available, and each operations can on average be processed on $1$ machine.
Figure 2: Explanation of the standard FJSSP benchmark specifications on the basis of the 11th test instance from the collection in [7].
The remaining rows of the text file then represent the individual jobs $J_{i}$ $(i=1,...,n)$ of the problem instance. The first value in each row gives the number of operations $n_{i}$ to be executed in job $J_{i}$. The second number is the number of machines $\lvert\mathcal{M}_{i,1}\rvert$ available to process the first operation $O_{i,1}$. This is followed by a sequence of $\lvert\mathcal{M}_{i,j}\rvert$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$. Next, the data for the second operation $O_{i2}$ are provided. In this way, all operations of all jobs are specified accordingly.
In order to facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance are given in the Appendix. In addition to the key properties such as the total number of jobs ($n$), machines ($m$), or operations(N) of an instance, seven other instance-specific statistics are provided that allow for a rough intuition of the individual benchmark instance complexity. These include the minimum ($T_{\min}$), maximum ($T_{\max}$) and average processing times ($\bar{T}$) measured over all possible machine-operation combinations, as well as the corresponding standard deviation $\sigma_{T}$. Moreover, the average number of operations per job $\frac{\texttt{N}}{n}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided. The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to.  [12]
The $\beta$-flexibility of a benchmark instance is determined as the ratio of the average amount of available machines per operation $m_{avg}$ and the total number of machines $m$ as
$\begin{split}\beta=\cfrac{m_{avg}}{m}\end{split}$ (2)
It indicates the degree of resilience in the FJSSP setting [13]. In case, all operations of all jobs can be processed by any machine, $\beta=1$ and the FJSSP is called fully flexible. Otherwise, i.e. $10<\beta<$ 1, it is only partially-flexible and, consequently, harder to solve. The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{ijk}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}\lvert\mathcal{M}_{i,j}\rvert}\end{split}$ (3)
Low $dv$ values indicate either very similar processing times across all available machines or a large amount of assignment options, respectively. This corresponds to many distinct scheduling solutions that have the same level of fitness.
Figure 3: Illustration of duration variety $dv$ and flexibility $\beta$ across all considered FJSSP instances.
Figure 3 displays scatter plots of two characteristic features of all benchmark instances. This provides an first overview of the diversity of the individual instances and reveals where similarities coincide. Details on the individual characteristics of the considered benchmark sources can be found in the Appendix. This rough insight into the benchmark properties alone supports the demand for a targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers.
2.3 Recent Work on Benchmarking Worker Flexibility in FJSSP
In order to retrieve benchmark instances including worker flexibility, the instances are usually specifically constructed for the given use case. In most cases, the constructed instances are based on subsets of commonly used benchmark instances [14][15][16]. In other cases, benchmark instances are randomly generated or constructed [16][17][18]. While the construction process is sometimes explained ([19][14]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems. The algorithm used to enable the conversion of all FJSSP benchmark instances was developed similarly to the algorithm used by [19].
[18] tries to balance the makespan, maximum worker workload, and total workload of machines using an multi objective discrete teaching-learning based optimization approach. They use 30 newly constructed benchmark instances. [14] minimize makespan, labor cost and green production related factors using a non-linear integer programming model, as well as a non-dominated ensemble fitness ranking algorithm. The 31 benchmark instances used, are constructed based on a subset of the commonly used FJSSP benchmark instances. [15] use a simulation based approach to optimize the FJSSP with worker flexibility and uncertain processing times with the minimum makespan and minimum standard deviation of makespan as objective. 36 benchmark instances are used for evaluation, constructed based on a subset of the commonly used FJSSP instances. [16] compare a mixed-integer linear programming approach with NSGA-II in a distributed FJSSP with worker flexibility on a subset of the common benchmarks (50 instances) and randomly generated instances (8 instances). The objectives for their optimization are makespan minimization, balanced machine workload and workload of workers. [17] considers the JSSP with worker flexibility on newly generated benchmark instances, focusing on the maximum worker workload, minimizing the makespan and balancing the human energy expenditure. To evaluate their approach, 40 problem instances are generated.
3 Representations of the Flexible Job Shop Scheduling Problem
This section establishes the foundation for setting up the benchmarking environment by describing three different modeling approaches for the FJSSP. The first of these is a Mixed Integer Linear Programming (MILP) representation defined in Sec. 3.1. It allows exact methods of mathematical programming to be applied. Furthermore, an implementation as Constraint Programming (CP) model is provided in Sec. 3.2. Toolboxes for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks. Lastly, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms, e.g. meta-heuristics, is presented in Sec. 3.3. It includes the definition of a precise schedule encoding and the design of the production system as an functional simulation environment.
3.1 Mixed-Integer Linear Programming Formulation
The Mixed Integer Linear Program (MILP) formulation is a proven model for the FJSSP to address these problem class using mathematical optimization methods, i.e. branch-and-bound or cutting-plane methods. In this section, the MILP is systematically formulated describing the decision variables, constraints and objective functions in detail.
Accordingly, the decision variables are defined in the following way:
Notation Description of decision variables Domain
$X_{i,j,i^{\prime},j^{\prime}}$ Indication of whether operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ on the same machine: $1$ if True, 00 else. $\mathds{Z}_{2}$
$Y_{i,j,k}$ The variable indicates whether operation $O_{i,j}$ is processed on machine $M_{k}$: $1$ if True, 00 else. $\mathds{Z}_{2}$
$C_{i,j}$ A variable representing the completion time of the $j$th operation $O_{i,j}$ in the $i$th job $J_{i}$ $\mathds{R}_{+}$
In this context, the notation $\mathds{Z}_{2}=\{0,1\}$ represents the ring of integers modulo $2$ indicating the binary character of a decision variable. Continuous variables, on the other hand, take on positive real numbered values in $\mathds{R}_{+}$. As a result, each schedule $\mathbf{y}\in\mathcal{Y}$ is represented by a feasible value assignment of these $(2+\texttt{N})\texttt{N}$ decision variables.
In addition, an index set $M_{i,j}$ is introduced that contains the indices $k$ of the permissible machines $M_{k}$ ($k=1,\dots,m$) for all operations $O_{i,j}$ ($j=1,\dots,n_{i}$) of all jobs $J_{i}$ ($i=1,\dots,n$):
$M_{i,j}=\{k:\text{$O_{i,j}$ can be performed on machine }M_{k}\}$ (4)
Next, the constraints of the FJSSP are described on the basis of these decision variables. To do so, the assignment, sequencing, and variable type constraints are considered independently. Afterwards, address a typical optimization objective associated constraint formulations.
3.1.1 General restrictions in FJSSP
Assignment constraints
It is imperative that the total number of $\texttt{N}=\sum_{i=1}^{n}n_{i}$ operations $O_{i,j}$ of an FJSSP instance be executed on exactly one of the machines available and admissible for the respective operation. These are specified by the indices in the set $M_{i,j}$.
[…]
By requiring that the sum of $Y_{i,j,k}$ over all admissible machine indices equals $1$, this set of N constraints guarantees that all operations $O_{i,j}$ are assigned to exactly one machine.
Sequencing constraints on the job level
The FJSSP demands that all operations of a single job need to be executed in a predefined sequence. For this type of restrictions, a novel formulation is proposed that involves the completion times $C_{i,j}$ of the operations within the FJSSP instance. Hence, one can require that the completion time $C_{i,1}$ of the first operation $O_{i,1}$ of job $i$ cannot be smaller than the corresponding processing time (see Eq. (6)). Also, the completion time of the $j$th operation $O_{i,j}$ of job $i$ cannot be smaller than the sum of the completion time of its predecessor operation $O_{i,j-1}$ in the same job $i$ and its own processing time (cf. Eq. (7)). Consequently, one receives another two sets of N constraints of the form
[…]
Sequencing constraints on machine level
Taking into account the order of operations on the machine level, it must be ensured that the completion times of two successor operations on the one and same machine are not unacceptably close to each other. For the proper formulation, an auxiliary large penalty parameter $L$ is introduced. This can, for example, be achieved by aggregating the maximal processing times of all operations over all machines
$L=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n_{i}}\max\nolimits_{k=1,\dots,m}(T_{i,j,k}).$ (8)
Accordingly, the following two sets of constraints are obtained
$C_{i,j}\geq$ $C_{i^{\prime},j^{\prime}}+T_{i,j,k}-L\cdot(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k}-Y_{i^{\prime},j^{\prime},k})$ (9)
$C_{i^{\prime},j^{\prime}}\geq$ $C_{i,j}+T_{i^{\prime},j^{\prime},k}-L\cdot(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k}-Y_{i^{\prime},j^{\prime},k}).$ (10)
Both sets (9) and (10) are required to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ as well as operations $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$ on those machines $M_{k}$ that satisfy $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$.
The two sets of constraints (9) and (10) are trivially fulfilled as long as the bracketed factor multiplied by $L$ is nonzero. In case that operation $O_{i,j}$ and $O_{i^{\prime},j^{\prime}}$ are processed on the same machine $M_{k}$, i.e. $Y_{i,j,k}=1$ and $Y_{i^{\prime},j^{\prime},k}=1$, the sequencing between these two specific operations is controlled by the decision variable $X_{i,j,i^{\prime},j^{\prime}}$. If $X_{j,k,j^{\prime},k^{\prime}}=1$ holds, operation $O_{i,j}$ is executed after operation $O_{i^{\prime},j^{\prime}}$ on the same machine $M_{k}$. This implies that the first set of constraints defined by Eq. (9) is active. It ensures that the processing time of the successor operation $O_{i,j}$ does not start before the predecessor operation $O_{i^{\prime},j^{\prime}}$ is finished. Conversely, if $O_{i^{\prime},j^{\prime}}$ is executed after operation $O_{i,j}$ on the same machine, the decision variable is $X_{i,j,i^{\prime},j^{\prime}}=0$. That is, the second set of constraints (10) is active and guarantees non overlapping processing times of both operations, respectively.
Variable type constraints
This type of constraints restrict the assignment of decision variables to permissible values. They are specified by the definition of the decision variables above. The binary variables $X_{i,j,i^{\prime},j^{\prime}}$ and $X_{i,j,k}$ only accept the values 00 and $1$ in $\mathds{Z}_{2}$. The range of the continuous variables $C_{i,j}$ is restricted to positive real values in $\mathds{R}_{+}$. In summary
$X_{i,j,i^{\prime},j^{\prime}}\in\{0,1\},\qquad Y_{i,j,k}\in\{0,1\},\quad\text{and}\quad C_{i,j}>0.$ (11)
Without considering a specific objective function $C$ and potentially associated additional constraints, the restrictions of the FJSSP environment are now adequately described by the constraints in Eqs. (5), (6), (7), (9), (10), and (11).
3.1.2 Objective function and associated constraints
Makespan minimization
The total time required for the execution of all jobs in an FJSSP instance is referred to as the makespan of the FJSSP. Its minimization represents one of the most common objectives considered in this problem domain. The formal formulation of the makespan depends on the representation of the FJSSP and in this specific case coincides with the maximum value $C_{max}$ of the completion times $C_{i,j}$ of all operations $O_{i,j}$. That is, the objective becomes
$\min\big{(}C_{\max}\big{)}=\min\left(\max_{\tiny\begin{array}[]{c}1\leq i\leq n\\ 1\leq j\leq n_{i}\end{array}}C_{i,j}\right)$ (12)
When using the above definition of the makespan, it must be guaranteed that the completion time of the last operation $O_{i,n_{i}}$ in each job $J_{i}$ is not exceeding the makespan. Consequently, one gets an additional set of constraints N associated with this specific objective
$C_{max}\geq C_{i,n_{i}}\qquad\text{ for }1\leq i\leq n.$ (13)
3.2 Constraint Programming Formulation
The FJSSP can be easily formulated as a discrete finite constraint optimization problem(COP). The two COP solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems.
In this comparison OR-Tools (ORT) an open-source solver developed by Google and CP-Optimizer (CPO) a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio are used. The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see 3.1).
3.2.1 Variables
The parameters representing the data are $T_{i,j,k}$ the processing time of operation $O_{i,j}$ on machine $k$. The variables are an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$ and an alternative interval variables $a\_I(i,j,k)$ for $k\in M_{i,j}$ to represent the feasible processing modes i.e. the selectable machines. For these intervals $a\_duration(i,j,k)=T_{i,j,k}$ holds. The ORT solver additionally uses a boolean variable $a\_presence(i,j,k)$ which indicates whether the alternative interval $a\_I(i,j,k)$ is chosen.
3.2.2 Objective
The objektive, i.e. minimizing the makespan, is formulated as
$\max_{1\leq i\leq n}(end\_of(I_{i,n_{i}}))\to\min$
3.2.3 Constraints
Assignment constraints
The assignment of an operation $O_{i,j}$ to a machine is done by linking the interval variable $I(i,j)$ to an alternative interval variable $a\_I(i,j,k)$.
This can be done by using the $alternative$ constraint of CPO. For $1\leq i\leq n$ and $1\leq j\leq n_{i}$ we add
$alternative\left(I(i,j),[a\_I(i,j,k):\mbox{ for }k\in M_{i,j}]\right).$
For the ORT solver the $exactly\_one$ constraint is used. For $1\leq i\leq n$ and $1\leq j\leq n_{i}$ we add
$exactly\_one$ $\left(a\_presence(i,j,k):\mbox{ for }k\in M_{i,j}\right)$
$start(i,j)$ $=a\_start(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
$duration(i,j)$ $=a\_duration(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
$end(i,j)$ $=a\_end(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
Sequencing constraints on job level
The sequencing constraints within jobs are formulated slightly different for the two solvers.
The CPO solver provides an $end\_before\_start$ constraint,
$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
and with the ORT solver starting and ending times are used
$start(i,j)\geq end(i,j-1)\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
Sequencing constraints on machine level
The $no\_overlap$ constraint which is provided by both solvers makes it very easy to formulate the sequencing constraint within machines. For $1\leq k\leq m$ the following constraints are added:
$no\_overlap\left(a\_I(i,j,k):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{if }k\in M_{i,j}\right)$
3.3 Simulation-Based Optimization Attempt
Multiple problem encodings can be found in the literature, the most common one being [19, 12]. To encode the information using this system, the FJSSP is split into 2 vectors. The first vector contains the sequencing information while the second vector contains the machine assignments. The sequence vector is encoded using the job IDs rather than the specific operation IDs. Since all operations can only be scheduled after the previous operation in the job sequence is completed, the specific operation can be determined by the number of times the job ID has previously appeared in the sequence vector. The assignment vector contains the IDs selected for the operation in a fixed order. For example, for a problem with 2 jobs, job 1 consisting of 3 operations and job 2 consisting of 2 operations, the sequence vector must contain the values $\mathbf{\{1,1,1,2,2\}}$ in any order. Assuming there are 2 machines in the production environment and the instance is fully flexible (every operation can be performed on any machine), the assignment vector can contain any combination of $1$ and $2$, with a length corresponding to the amount of operations. The order of the values is fixed, starting from the first operation of the first job, moving to the last operation to the first job, then moving on to the first operation of the second job and moving through to the end to the last operation of the second job. One example for the described problem could look like Listing 1.
1s = [1, 2, 1, 1, 2]
2a = [2, 2, 1, 2, 1]
Listing 1: Example sequence and assignment vector
$\mathbf{s}=\begin{pmatrix}1&2&1&1&2\end{pmatrix}^{\top}\qquad\textrm{and}\qquad\mathbf{a}=\begin{pmatrix}2&2&1&2&1\end{pmatrix}^{\top}$ (14)
The resulting schedule is depicted in Fig. 4.
Figure 4: Example schedule produced by Listing 1.
Each schedule $\mathbf{y}=(\mathbf{s}|\mathbf{a})\in\mathcal{Y}\subset\mathds{N}^{\texttt{N}+\texttt{N}}$ is encoded by two distinct vectors $\mathbf{s}$, and $\mathbf{a}$. The sequence vector $\mathbf{s}\in\mathds{N}^{\texttt{N}}$ refers to the job sequence of the schedule and determines the order of the operations on their assigned machines. A feasible $\mathbf{s}$ contains each job index $i$ exactly $n_{i}$ times and the $j$th appearance of index $i$ corresponds to operation $O_{ij}$. Since the sequence of operations of a job is fixed, the job dependent sequence will always be feasible. The assignment vector $\mathbf{a}\in\mathds{N}^{\texttt{N}}$ contains the corresponding machine assignments. The order in $\mathbf{a}$ is fixed in the beginning and never changes to make sure each slot can only be assigned with available machines for every operation. This, together with the encoding chosen for $\mathbf{s}$ ensures that all solutions produced either randomly or through recombination and mutation will always be feasible. To evaluate the solution, the schedule is constructed according to the instructions in $\mathbf{s}$ and $\mathbf{a}$. To construct the schedule, the operations are inserted according to the sequence in $\mathbf{s}$. The start time $t_{ij}$ is determined based on the current last operation ending time on the assigned machine $t_{a_{ij}}$ and the last ending time in the job sequence of the operation $t_{prev}$ (Eq. (15)).
$t_{prev}=\begin{cases}t_{i,j-1}+d_{i,j-1}&\text{if $j>0$}\\ 0&\text{if $j=0$}\end{cases}$ (15)
$t_{ij}$ is then determined by Eq. (16).
$t_{i,j}=\max(t_{prev},t_{a_{i,j}}).$ (16)
$t_{a_{,ij}}$ is replaced (Eq. (17)) and the next operation is inserted into the schedule.
$t_{a_{i,j}}=t_{i,j}+d_{i,j}.$ (17)
After all starting times have been determined, the solutions fitness $C_{max}$ is determined by Eq. (18).
$C_{\max}=\max_{\tiny\begin{array}[]{c}1\leq i\leq n\\ 1\leq j\leq n_{i}\end{array}}(t_{i,j}+d_{i,j})$ (18)
4 Expansion of the FJSSP for Worker Flexibility
The Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W) is an extension to the FJSSP which adds worker flexibility. This means that for every operation, an eligible worker also needs to be assigned to operate the machine to which the operation is assigned. The selected machine-worker combination further influences the processing time of the operation on the machine. In addition to the FJSSP, the FJSSP-W therefore introduces a set of workers $\mathcal{W}=\{W_{1},\dots,W_{o}\}$. Each operation $O_{i,j}$ needs to be assigned to a permissible machine $M_{k}\in\mathcal{M}_{i,j}\subset\mathcal{M}$ and a worker $W_{s}\in\mathcal{W}_{i,j,k}\subset\mathcal{W}$ the set of permissible workers to process the operation $O_{i,j}$ on the assigned machine $M_{k}$. Similar to the FJSSP, the FJSSP-W assumes that all operations, machines, and workers are operational and available at time 0. The objective for the FJSSP-W is to find those operation-machine-worker combination that minimize the makespan $C$. As opposed to the FJSSP, the flexibility $\beta$ of a benchmark instance needs to take more into account than only the machines. All possible assignment combination for every operation need to be taken into account. This is the same as for the FJSSP, but includes the possible worker assignments for each operation-machine combination as well. The average amount of assignment options $\omega_{avg}$ is determined by Eq. (20).
$\omega_{all}=\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)$ (19)
$\omega_{avg}=\frac{1}{\texttt{N}}\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\sum_{k=1}^{m}|\mathcal{W}_{ijk}|\quad$ (20)
Which leads to $\beta$ defined as Eq. (21), with $\omega_{uni}$ being the overall amount of unique machine-worker combinations available for the benchmark instance.
$\beta=\cfrac{\omega_{avg}}{\omega_{uni}}$ (21)
In the same vein, the definition for $dv$ for FJSSP-W needs to be rewritten as Eq. (22)
$\begin{split}dv=\cfrac{d_{unique}}{\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)}\end{split}$ (22)
4.1 Design Proposal of Worker Flexibility Integration
The FJSSP-W considers that operations in the production process need workers to be present. Therefore, workers also need to be assigned to the different operations. In the FJSSP-W, it is considered that not all workers will work at the same speed, and not all workers are capable of performing every operation with ever (or any) of the available machines for the operation. To receive a set of benchmark problems to test and develop appropriate approaches, the given FJSSP benchmark instances are extended to contain the additional required information about the possible worker assignments. To that end, we provide an algorithm similar to [20] , which is capable of extending any FJSSP instance in the format of the gathered benchmark instances (Alg. 1). The parameters used for the benchmark generation are the amount of workers (default: $1.5*m$), the upper and lower bound range for the durations (default: $ub=1.1$, $lb=0.9$ of the original duration) to have a random fluctuation in the durations required by the workers of ±$\pm 10\%$, and optionally a random seed for reproducibility.
After the initialization of the required parameters, in line 1 an empty list for saving the result benchmark is created. In line 2, the original benchmark file, located at $path$, is read and saved into a list, separated into the rows of the file. The base information about the benchmarks are added to the result, as a list representing a new row in the output, in line 3, keeping the same format as the original files, with the added information about the amount of workers available for the problem instance. Starting from line 4 and ranging to line 25, the rest of the rows of the original benchmark file are processed one by one. In line 5, a variable to keep track of the reading position in the current row ($idx$) is initialized, as well as an empty list, which is added to the output as a new row. The amount of operations needed for the job represented by the row is added to the new row in line 6. From line 7 to line 24, the current row from the original benchmark, according to the original format of the problem instances, is processed. To do so, $idx$ is raised by $1$ in row 8 to address the next operation that needs to be parsed. In line 9, the amount of machines available for the operation is read, and added to the output row in line 10. From line 11 to line 23, the different machine assignments and their respective durations for the current operation are parsed. To achieve that, line 12 raises $idx$ by $1$ again to advance the reading position of the input current row. Since $idx$ is now referencing to a machine ID in the original benchmark, this possible machine assignment is also added to the output row in line 13. In line 14, the amount of workers which are going to be able to perform the required operation on this specific machine is chosen uniformly from the interval [$1$, $n_{workers}$]. The amount of workers available for this machine assignment ($workers$) is added to the output row in line 15. In line 16, the available workers for the assignment are selected randomly, using a function named $Sample$, using the parameters $n_{workers}$ and $workers$ to select the IDs of the workers in the interval [0, $n_{workers}$). Note that generally the assigned IDs of machines and workers are converted to 0-based indexing for easier implementation as opposed to the original 1-based indexing. The subset of workers is stored in the list $\mathbf{selected}$. To advance the reading position again, $idx$ is raised by $1$ in line 17. $idx$ now points at the original duration of the machine assignment for the given operation, which is stored in $d_{orig}$ in line 18. Starting from line 19 and ending in line 22, the current operation-machine assignments are finalized by choosing the durations for the different worker assignments. To do this, the duration for each worker $d_{work}$ is uniformly chosen from the interval [$d_{orig}*lb$, $d_{orig}*ub$] and rounded to the closest integer in line 20. Line 21 adds the worker and the duration for the operation-machine-worker combination to the output row in $out$.
Algorithm 1 Adding Worker Flexibility to common FJSSP Benchmark instances
1:$n_{workers}>0$, $lb\geq 0$, $ub\geq 0$, $path$
2:$\mathbf{out}\leftarrow\{\}$
3:$\mathbf{rows}\leftarrow ReadFile(path)$
[…]
5:for $i\leftarrow 1$ to $||\mathbf{rows}||$ do
6:     $\mathbf{out}_{i}\leftarrow\{\}$, $idx\leftarrow 0$
7:     $\mathbf{out}_{i}\cup\mathbf{rows}_{i,idx}$
8:     for $j\leftarrow 1$ to $||\mathbf{rows}_{i}||$ do
9:         $idx\leftarrow idx+1$
10:         $m\leftarrow\mathbf{rows}_{i,idx}$
11:         $\mathbf{out}_{i}\cup\{m\}$
12:         for $k\leftarrow 1$ to $m$ do
13:              $idx\leftarrow idx+1$
14:              $\mathbf{out}_{i}\cup{\mathbf{rows}_{i,idx}}$
15:              $workers\leftarrow Uniform(1,n_{workers})$
16:              $\mathbf{out}_{i}\cup\{workers\}$
17:              $\mathbf{selected}\leftarrow Sample(n_{workers},workers$)
18:              $idx\leftarrow idx+1$
19:              $d_{orig}\leftarrow\mathbf{rows}_{i,idx}$
20:              for $l\leftarrow 1$ to $workers$ do
21:                  $d_{work}\leftarrow Round(Uniform(d_{orig}*lb,d_{orig}*ub))$
22:                  $\mathbf{out}_{i}\cup\{selected_{l},d_{work}\}$
23:              end for
24:         end for
25:     end for
26:end for
After this routine is finished, the benchmark is extended with worker flexibility and kept similar to the original format. The original format can be seen in an arbitrarily constructed example in Listing 2.
12 3
23 2 0 10 2 15 3 0 10 1 5 2 10 2 0 10 2 15
32 1 1 15 2 1 10 2 15
Listing 2: Original Benchmark example
Row 1 of Listing 2 contains the information that the problem instance consists of 2 jobs and 3 available machines. In row 2 it is specified that Job 1 consists of 3 operations. Operation 1 of Job 1 can be processed on 2 machines (Machine 0 with processing time 10 and Machine 1 with processing time 15). Operation 2 of Job 1 can be processed on 3 machines. Either Machine 0 with processing time 10, Machine 1 with processing time 5 or Machine 2 with processing time 10. Operation 3 of Job 1 has 2 machines available (Machine 0 with processing time 10 and Machine 2 with processing time 15). Job 2 requires 2 operations. Operation 1 of Job 2 can be processed on only 1 machine (Machine 1 with processing time 15), while Operation 2 of Job 2 can be processed on 2 machines (Machine 1 with processing time 10 and Machine 2 with processing time 15).
In the new format of the benchmarks, line 1 contains the general information about the instance (amount of jobs, amount of machines, amount of workers), followed by a row by row description of every job that needs to be scheduled. Every row starts with the number of operations required for the job to be finished. Following this first row entry, every operation starts with the amount of machines required, followed by the entries for every machine. Every entry for the machine starts with the machine id, followed by the amount of workers available to process the operation on this specific machine. After that, for every worker available on the machine, an entry consisting of the worker id, followed by the duration of the operation-machine-worker combination is listed in the row.
An arbitrarily constructed example for this format can be seen in Listing 3. REPLACE:
1 2 3 4
2 3 2 0 2 1 10 3 5 2 1 0 15 3 0 1 2 10 1 2 0 10 3 5 2 2 2 5 3 15 2 0 2 2 15 3 10 2 2 0 15 1 5
3 2 1 1 3 1 10 2 5 3 10 2 1 1 2 10 2 2 0 10 3 15
Listing 3: Benchmark with worker flexibility example
This example means that, according to row 1, there are 2 jobs, 3 machines and 4 workers in this production environment. Job 1 (row 2) consists of 3 operations. Operation 1 can be processed on 2 machines. The first of these machines is Machine 0, which has 2 workers available to perform this operation. Worker 1 with a processing time 10 and Worker 3 with a processing time of 5. The second machine to process Operation 1 of Job 1 is Machine 2. Machine 2 has 1 worker available (Worker 0) with a processing time 15. Operation 2 of Job 1 has 3 machines available. Machine 0, with 1 worker (Worker 2) with processing time 10, Machine 1 with 2 workers (Worker 0 with processing time 10 and Worker 3 with processing time 5), and Machine 2 with 2 workers (Worker 2 with processing time 5 and Worker 3 with processing time 15). Operation 3 of Job 1 can be processed on 2 machines. Machine 0, which has 2 possible worker assignments (Worker 2 with processing time 15 and Worker 3 with processing time 10), and Machine 2. Machine 2 also has 2 possible workers (Worker 0 with processing time 15 and Worker 1 with processing time 5). Job 2 (row 3), requires 2 operations to be processed. Operation 1 of Job 2 has 1 possible machine (Machine 1) with 3 possible workers. Worker 1 with processing time 10, Worker 2 with processing time 5 and Worker 3 with processing time 10. Operation 2 of Job 2 can be assigned to 2 machines. Either Machine 1, which has 1 possible worker (Worker 2 with processing time 10), or Machine 2, with the possibility to assign the operation to 2 different workers (Worker 0 with processing time 10 or Worker 3 with processing time 15).
4.2 Mixed-Integer Linear Programming Formulation
An obvious generalization of the MILP formulation for the FJSSP (3.1) is used. The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in 3.1. The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
4.2.1 Indices and Variables
To denote the workers the subscripts $s,s^{\prime}$ with $1\leq s,s^{\prime}\leq w$ and $w$ the number of workers are used. The processing time of operation $O_{i,j}$ on machine $M_{k}$ also depends on worker $W_{s}$ and is denoted by $T_{i,j,k,s}$
The decision variables $Y_{i,j,k}$ in 3.1 have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else. Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker. To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
$W_{i,j,k}=\{s:\text{ such that $O_{i,j}$ can be processed on $M_{k}$ by worker$W_{s}$}\}$
is used.
4.2.2 Constraints
Assignment constraints
These constraints guarantee that all operations $O_{i,j}$ are assigned to exactly one machine and one worker.
$\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}=1\text{ for }1\leq i\leq n,1\leq j\leq n_{i}$
Sequencing constraints on job level
$C_{i,1}\geq\sum\nolimits_{k\in M_{i,1}}\sum\nolimits_{s\in W_{i,1,k}}Y_{i,1,k,s}T_{i,1,k,s}\text{ for }1\leq i\leq n$
$C_{i,j}\geq C_{i,j-1}+\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}T_{i,j,k,s}\text{ for }1\leq i\leq n,1<j\leq n_{i}$
Sequencing constraints on machine level
$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k,s^{\prime}}-L(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
Both constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$, for machines $M_{k}$ with $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$ and for workers $W_{s}$ ($s\in W_{i,j,k}$) and $W_{s^{\prime}}$ ($s^{\prime}\in W_{i^{\prime},j^{\prime},k}$).
Sequencing constraints on worker level
$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-U_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k^{\prime},s}-L(U_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
As in 4.2.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$. But contrary to 4.2.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$. For the large number $L$, one can use
$L=\sum\nolimits_{i=1}^{n}\sum\nolimits_{j=1}^{n_{i}}\max\nolimits_{1\leq k\leq m;1\leq s\leq w}(T_{i,j,k,s})$
The idea of these constraints is the same as in Eq. (9) and (10) explained there in detail. Finally makespan and variable type constraint are the same as in 3.1.
4.3 Constraint Programming Formulation
The same as for the MILP formulation holds also for the CP formulation: The FJSSP-W is an obvious generalization of the FJSSP. In the following the formulation for CPO solver is presented. The necessary changes for using the ORT solver are obvious when the two different formulations for the FJSSP are kept in mind. Again, each operation is represented by an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$. For the FJSSP-W an alternative interval variable $a\_I(i,j,k,s)$ has to be defined for each admissible machine-worker combination.
The assignment of an operation $O_{j,k}$ to exactly one machine-worker combination is done by linking the interval variable $I(j,k)$ to an alternative interval variable $a\_I(i,j,k,s)$. This can be done by the $alternative$ constraints
$alternative\left(I(i,j),[a\_I(i,j,k,s):\mbox{ for }k\in M_{i,j},s\in W_{i,j,k}]\right)$
for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
The sequencing constraints on job level are the same as in 3.2:
$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
and the sequencing constraints on machine level are very similar to 3.2. For $1\leq k\leq m$ the following constraints are added:
$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{ if }k\in M_{i,j}\mbox{ and }s\in W_{i,j,k}\right)$
Additionally the sequencing on worker level has to be guaranteed by using another set of $no\_overlap$ constraints. For $1\leq s\leq w$:
$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i},k\in M_{i,j}\mbox{ if }s\in W_{i,j,k}\right)$
4.4 Simulation-Based Optimization Attempt
The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section 3.3. To add the worker flexibility, an additional vector is added. The vector has the same properties as the machine assignment vector, in that it is in the same fixed order and contains the IDs of the workers assigned to the operation-machine combination chosen to do the task. Extending the example used in section 3.3 with 3 available workers for the system, a problem instance could look like Listing 4.
1s = [1, 2, 1, 1, 2]
2a = [2, 2, 1, 2, 1]
3w = [3, 1, 1, 2, 3]
Listing 4: Example sequence, assignment and worker vector
The resulting schedule looks as depicted in Fig. 5.
Figure 5: Example schedule produced by Listing 1.
5 Description of the Benchmarking Environment
5.1 Implementation details
The benchmarking environment is implemented in Python [21]. Several APIs are offered by the environment, including the evaluation of FJSSP and FJSSP-W solutions and their comparison to the known results.
Additionally, the benchmarking environment contains the best known values for the FJSSP and FJSSP-W instances created with the same parameters than the instances used in this paper. The algorithm which turns FJSSP instances into FJSSP-W instances is also provided.
5.2 Performance evaluation
For comparisons to other solvers the benchmark environment offers different visualizations of the results over all used benchmark instances. The data of the evaluated solvers are included and best known results can be updated over time.
(a)Refer to caption (b)Refer to caption
Figure 6: An example plot using example data displaying the portion of instances that lie within a certain gap from the best found solution. A gap of 1.0 corresponds to 100% distance. In (a), all of the data is displayed, while (b) shows an example with a more zoomed in perspective on the same data (limit on the x-axis as parameter to the visualizer code)
The example plot shown in Fig. 6(a) compares the relative deviation (gap) $\delta_{rel}$ (Eq. (23)) of the best solution of a solver $C_{fb}=C(\mathbf{y}_{fb})$ for each algorithm from the best-known result $C_{best}$. On the y-axis, the plot shows the portion of instances where the solutions were within the given gap to $C_{best}$.
$\begin{split}\delta_{rel}=\frac{C_{fb}-C_{best}}{C_{best}}.\end{split}$ (23)
Fig. 6(b) shows the same (on example data) with a zoomed in perspective (limit on the x-axis).
(a)Refer to caption (b)Refer to caption
Figure 7: Based on example data, the portion of problem instances per solver that were optimized within 10% of the best found result. (a) shows the results in relation to the best solution of all solvers within, while (b) shows the time needed for the instances to reach the threshold in relation to their own best solution.
The plots shown in Fig. 7 show the portion of instances which reached the given criteria in $n$ seconds. The example plots use the fitness as example, but the same plots can also be done for other criteria, e.g. the lower bounds found by a solver. The maximal distance to the target value can be given to the visualizer as a parameter. In the case of Fig. 7(a), the plot shows the portion of instances capable of getting within a range of no more than 10% distance from the overall best known result $C_{best}$ within the given time frames. Fig. 7(b) however shows how long it took each solver to get within the range of 10% of it’s own best solution $C_{fb}$ for which portion of the test instances. Note that while for Fig. 6(a) all solvers reach 100% eventually, granted that all instances included in the comparison have been solved by each solver, in Fig. 7(a) there may be many instances that never reached the threshold. The visualizer is implemented in a way that it can handle missing data for the case of Fig. 6(a). The corresponding line for the solver with the missing results would not reach 100%, which is the only way this can happen for these plots (without a limit imposed on the x-axis).
6 Baseline Experiments
6.1 Experimental Settings
For the FJSSP-W, the solvers were tasked with solving all 402 benchmarks. The hardware used for the experiments is a Intel Core i7-6700 (3.40GHz) CPU with 16 GB RAM. All solvers were given a maximum of 1200 seconds (20 minutes) for each benchmark instance. The benchmark instances, including all results are available online [22].
6.2 Applied Solvers
All used solvers, including the self implemented GA, were not manually tuned or configured to accommodate any of the benchmark instances and were run in their default settings. However, some of the solvers do automated pre-processing and parameterizing on their own, e.g. Gurobi or the self implemented GA, which determines some of its parameters automatically based on the benchmark characteristics.
Hexaly
Hexaly is a new kind of global optimization solver. It offers nonlinear and set-oriented modeling APIs. Relying on innovative proprietary algorithms, it is faster and more scalable than Mixed-Integer Programming, Constraint Programming, and Nonlinear Programming solvers. [23]
OR-Tools
OR-Tools is an open-source software suite for optimization [24] developed by Google. It offers several ways to solve constraint programming and linear programming problems. The solver used for the results in this paper is the CP-SAT solver, which is a constraint programming solver that uses SAT (satisfiability) methods.
ILOG CPLEX
CPLEX by IBM is an optimization software package used to solve several optimization problem types, including integer programming, linear programming, quadratic programming problems, and constraint programming problems. CPLEX is used for both its constraint programming capabilities (CPLEX-CP) and its MILP solver capabilities (CPLEX-LP). [25]
Gurobi
Gurobi is a commercial mathematical solver commonly used to solve MILP problems and is developed by Gurobi Optimization LLC. [26]
Genetic Algorithm
To solve the FJSSP-W, an extended version of the GA implemented in [12] is used. Additionally, the implementation is done in the programming language C# rather than Python. To evaluate the GA properly, all instance are repeated 18 times to account for the randomness of the GA. The 18 repetitions are a result of the employed hardware being able to handle 6 instances of the GA running in parallel processes. With a maximum running time of 20 minutes, all repetitions for all instances were expected to take 402 hours (16.75 days). However, this runtime was reduced by preemptively stopping the GA once it reached the optimal result. However, other than the CP and MILP solvers, the GA has no built-in method to determine if a found solution is optimal. Therefore, the GA was run as the last solver and relied on results reported by the other solvers as optimal results, if any optimal result was found for a problem instance. This helps illustrate which ranges of the different characteristics are easiest to handle. For example, it can be observed that the difficulty of a problem increases with a higher number of operations, as well as with a lower duration variety in the processing times.
6.3 Baseline Results
Figure 8: CDF plot showing selected characteristics of all problem instances
The tables for each solver show the number of instances in which the solver reached the best result, the average characteristics of the instances and the respective standard deviations. The rows in the tables separate the instances in which the solver was the only solver to reach the best result (Single Best), the instances in which at least one other solver also reached the best result (Shared Best), an overview of all instances for which the solver found the best result (Combined), which includes the 2 prior rows, and lastly the instances for which the solver reached the optimal solution (Optimal). Note that since the GA has no way to confirm if a solution is an optimum, the best results for each instances were collected. For each solver the instance counts if it reached the known optimum even if it did not identify it as the optimum, as long as one of the solvers confirms the result to be optimal. Fig. 8 shows the cumulative distribution function over all used benchmark instances on selected problem characteristics. The vertical lines mark the average characteristic for instances where the best result was achieved by one solver (One Best), by a subset of the solvers (Multiple Best), and instances where all solvers could find the same best result (not necessarily the optimal result).
Table 2: Characteristics of all benchmarks.
mean std max min
$\beta$ 0.2806 0.1749 1.0 0.0667
$dv$ 0.1496 0.2241 0.95 5.6669e-05
N 150.1144 111.2573 500 4
$m$ 13.5124 13.1667 60 2
6.3.1 MILP
Gurobi
The characteristics of the instances Gurobi found the best solutions for are shown in Tab.3. Gurobi could not solve any instance that could not also be solved by at least one other solver. The instances that can be solved by Gurobi all reach the known optimum (11). The instances in question are small in terms of number of machines and number of operations. The flexibility is high in comparison to the average flexibility in all benchmarks (Tab. 2). The same is true for the duration variety.
Table 3: Characteristics of benchmarks in which Gurobi achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 0 - - - - - - - -
Shared Best 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
Combined 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
Optimal 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
CPLEX-LP
The characteristics of the instances the LP solver included in CPLEX found the best solutions for are shown in Tab. 4. The attributes are similar to the instances solved by Gurobi. While CPLEX-LP also could not solve any instances as the only solver, it managed to solve 11 instances to the known optimum. The instances solved by CPLEX-LP are smaller than the ones solved by Gurobi on average. The flexibility is slightly lower, while the duration variety is significantly higher.
Table 4: Characteristics of benchmarks in which the LP solver included in CPLEX achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 0 - - - - - - - -
Shared Best 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
Combined 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
Optimal 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
6.3.2 CP
OR-Tools
The characteristics of the instances OR-Tools found the best solutions for are shown in Tab. 5. OR-Tools managed to reach the best solution in 15 instances that could not be solved as well by any other solver. It also reached the best result for 56 instances for which at least one other solver reached the best result. Of the 71 best results reached, 48 are confirmed to be optimal. The average amount of machines is slightly higher than for both MLIP approaches, while the standard deviation is similar. Compared to CPLEX-CP and the GA the amount and spread of machines is low, but similar to Hexaly. For the amount of operations, OR-Tools solves bigger instances than both MILP approaches, but smaller than Hexaly, CPLEX-CP and the GA. The flexibility is smaller than the MILP approaches, CPLEX-CP, and GA, and similar to Hexaly. In comparison to Hexaly, CPLEX-CP and the GA, OR-Tools requires a higher duration variety to succeed, but can handle a lower duration variety than both MILP approaches, considering that the duration variety does not significantly change for the different categories shown in the table.
Table 5: Characteristics of benchmarks in which OR-Tools achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 15 7.47 2.45 78.0 29.33 0.11 0.06 0.3061 0.2607
Shared Best 56 6.11 2.96 53.02 40.21 0.39 0.25 0.3912 0.302
Combined 71 6.39 2.90 58.29 39.35 0.36 0.24 0.3732 0.294
Optimal 48 5.67 2.23 47.31 33.58 0.43 0.26 0.3960 0.296
CPLEX-CP
The characteristics of the instances CPLEX found the best solutions for are shown in Tab. 6. CPLEX-CP managed to find the best solution in 283 cases where no other solver found the best solution. The best result was also found for 66 additional benchmark instances for which other solvers also found the best solution. 69 of the overall 349 found best results are the known optimum of that benchmark instance. CPLEX-CP can handle lower duration varieties than the MILP approaches and OR-Tools, as well as slightly lower duration varieties than Hexaly. However, it can not handle the lowest duration varieties like the GA. In terms of the flexibility, the average flexibility of the instances only CPLEX-CP could solve best is slightly lower than the flexibility where the GA performs best, but higher than for all the other solvers. The average values for the instances that only CPLEX-CP solved best are very close to the average instance attributes (Tab. 2) of all benchmark instances. CPLEX-CP can handle a high operation amount in comparison to all other solvers except the GA. The average amount of machines is also higher than for the other solvers, the exception again being the GA.
Table 6: Characteristics of benchmarks in which CPLEX achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 283 14.14 12.55 175.96 112.49 0.28 0.15 0.1026 0.177
Shared Best 55 15.99 21.27 63.53 51.41 0.04 0.06 0.3424 0.294
Combined 338 14.20 13.89 158.14 114.98 0.014 0.03 0.1361 0.216
Optimal 69 20.77 21.83 73.99 62.06 0.039 0.05 0.3059 0.303
6.3.3 SBO
The characteristics of the instances the GA found the best solutions for are shown in Tab. 7. The GA could find the best solution for 31 instances, 17 of which are the known optimum. Any of the 7 best solutions that only the GA could find could also be the optimum, but the GA is lacking any method of confirming the optimum. Note that the stable value of 60 for amount of machines point at the Behnke instances [5], which contain the largest problem instances. For the instances that only the GA could solve best, the amount of machines and operations are high on average with a big range especially for the amount of operations. Paired with the flexibility values, which are slightly higher than CPLEX-CP and significantly higher than for the other solvers, this indicates that the GA is well suited for large and complex instances. Additionally, the GA performed better than all other solvers for instances with a low duration variety, which makes these instances harder to solve. Through the average metrics for the Shared Best instances, it can be seen that the GA can still also perform well on instances with high duration variety, and higher flexibility. Measuring the stability of the GA for every benchmark instance using Eq. (24), the average spread around the average result over all instances is 6.35% (median 5.01%).
$\begin{split}\delta=\cfrac{f_{worst}-f_{best}}{f_{average}}\end{split}$ (24)
For the instances for which the GA could find the best result overall, the average amount of machines stands out compared to all other solvers. The average amount of operations is also higher than the other solvers, though Hexaly comes close to a similar problem size, but with a much lower standard deviation. The average flexibility is below OR-Tools but above CPLEX-CP.
Table 7: Characteristics of benchmarks in which the GA achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 7 60.00 0.00 214.29 207.59 0.16 0.02 0.0003 0.0003
Shared Best 24 11.75 16.73 44.33 71.21 0.34 0.21 0.4393 0.358
Combined 31 22.65 25.20 82.71 133.13 0.29 0.19 0.3402 0.3648
Optimal 17 4.65 2.42 12.12 7.12 0.39 0.19 0.6181 0.2603
The characteristics of the instances for which Hexaly found the best solutions are shown in Tab. 8. Overall, Hexaly found the best solution for 71 of the instances, 37 of which were identified as the optimum (by any of the solvers). 29 of the best results were on instances where only Hexaly could find the best result, and 42 of the best results were also found by at least one other solver. Hexaly performs well on instances with a higher duration variety than the GA and CPLEX-CP but handles a lower $dv$ better than OR-Tools and the MILP approaches. Hexaly seems to deal best with the lowest flexibilities out of all solvers. The problem size is narrower spread than for the Single Best results of the GA and CPLEX-CP and on average lower, but higher than the other solvers. The same applies to the amount of machines.
Table 8: Characteristics of benchmarks in which Hexaly achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 29 10.10 1.92 116.76 52.25 0.07 0.05 0.1424 0.153
Shared Best 42 6.12 3.19 53.12 52.19 0.24 0.19 0.3887 0.2943
Combined 71 7.75 3.36 79.11 60.66 0.17 0.17 0.2881 0.2738
Optimal 37 5.43 2.15 50.27 54.41 0.26 0.19 0.3984 0.289
6.4 Comparison
Figure 9: The portion of problem instances per solver that are within a relative gap (Eq. (23)) to the best known result.
Optimal Feasible Infeasible
CPLEX_CP 69 331 2
CLPLEX_LP 11 310 81
Gurobi 11 305 86
Hexaly 37 365 0
OR-Tools 48 349 5
GA_Best 17 385 0
GA_Average 14 388 0
Table 9: Caption
The results in Fig. 9 show the relative gap of the results of each solver to the best result found between all solvers. While the results in [12] showed that MILP solvers perform better on the FJSSP instances, the results indicated that the GA would start to outperform the MILP solver on more complex problems. This can be shown through the data collected using the FJSSP-W instances. Considering the best results of the GA as well as the average results of the GA perform better than both used MILP solvers for the FJSSP-W. This shows that the GA can scale to more complex scenarios better. Fig. 9 shows that except for the MILP approaches, all solvers are less than 25% worse than the best overall found result on over 80% of all instances.
Given even more complex scenarios, like the FJSSP-W with added uncertainties for robust optimization, the GA can be expected to increase it’s advantage over the other solvers even further.
As seen in the results (see Sec. 6.3), all solvers show tendencies towards a certain category of problem instance. While the GA seems to scale best to large, complex problems, the CP solvers are closest to solving the average benchmark instance for the given benchmark collection. The known optima found by the solvers all have a higher than average duration variety, showing that a high duration variety indicates instances which are easier to solve across all solvers. The same seems to be true for the flexibility. Considering the combination of high duration variety and high flexibility, this means that for those instances the assignment problem can be solved easier, leaving the main pressure on solving the sequencing problem. However, the average amount of machines and operations on instances that are solved optimally is lower than the average benchmark instance, meaning that mostly small instances with high duration variety are the ones that can be solved by any of the solvers easily.
Table 10: Average characteristics of optimally solved benchmarks.
mean std
$\beta$ 0.4145 0.2251
$dv$ 0.03872 0.0542
N 73.9855 62.0568
$m$ 7.1739 3.3429
$d_{mean}$ 100 131.3469
$d_{stdv}$ 52.9665 79.3247
$d_{span}$ 189 272.8644
Figure 10: Benchmark attributes of all instances normalized over the maximum value.
The characteristics of the benchmarks that could be solved optimally are shown in Tab. 10. From this data it can also be seen that the solvers generally seem to profit from an easier assignment problem, with a low number of machines on average and a mid-range flexibility. For these smaller instances, a smaller duration variety also appears to be less of an issue, for the same reason. The distribution of actual duration values was also considered (Fig. 10) , since some problem instances deal with duration values smaller than 10 while other instances have duration values up to over 1000, with he maximum span present in the benchmark set being over 1600. However, due to the large variation of duration value metrics in the optimally solved instances (Tab. 10), the actual values do not seem to impact the solvers, as long as the duration variety decreases the number of solutions with the same fitness values and the amount of mutations applied to solutions that do not lead to change in their fitness value. In comparison to the average characteristics of all benchmarks (Tab. 2), the flexibility of the optimally solved instances is higher and the problem size is lower.
Overall the data implies that the assignment part of the optimization (machines for FJSSP, machines and workers for FJSSP-W) seems more problematic for all solvers but the GA, which additionally explains why the GA scales up better to the FJSSP-W than other solvers.
For further investigation, it could prove useful to craft a better balanced benchmark collection, varying in problem size (machines, workers, operations) as well as their complexity (flexibility, duration variety), to cover the metrics more evenly for testing and comparison purposes. Many of the different benchmark sources also contain multiple instances that cover very similar problems, which results in a lot of experiment runtime is used to test the solvers on the essentially the same problem.
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 11: Best solver for benchmark characteristics
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 12: Optimal solution found by the different solvers
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 13: Best solution based on solver category. While MILP approaches could not achieve the single best result for any instance, they could optimize 11 instances the same as all other solvers. CP approaches achieved by far the most singled out best results (315), followed by Hexaly (29), then the GA (8). 39 instances were solved best by a mixture of solver categories and 11 instances were solved the same by all solver categories.
7 Conclusion
While the GA only outperforms the MILP approaches on the FJSSP on specific instances  [12], the MILP approaches are not capable of scaling up to the FJSSP-W as well as the GA. Except for CPLEX-CP, the same is true for the other approaches, like the CP approach using OR-Tools and Hexaly. While both can still keep up with the GA on certain instances, the GA shows better potential in scaling up to more complex scenarios. In direct comparison between the CP approach and the GA approach, the characteristics of the benchmark problems solved best by either show some significant differences. Namely, the GA deals better with a more difficult assignment problem, with both more machine and workers as well as more operations. However, for simpler assignment problems (low $dv$), the CP approach outperforms the GA in runtime. While it is likely that the efficiency of the implementation plays a role for those problems, the CP approach does tend to find its final solution earlier on in the optimization process than the GA, which steadily increases its best solution over time. This shows that both approaches have an advantage in different scenarios, which can possibly be determined by the problem characteristics, and could be exploited by a hybrid model. The capability of the GA to handle low $dv$ values in large problem instances demonstrates its suitability for even more complex use-cases, including FJSSP and FJSSP-W with uncertainties.
8 Outlook
To make the current system more applicable to the real-world, the developed solvers need to be more robust towards uncertainties. Given that the GA shows stability when scaled to more complex problems, it is well geared to include uncertainty in its optimization process, while other approaches like MILP and CP struggle with uncertainty. However, given the efficiency of CP to find good solutions, a hybrid method combining the advantages of both CP and GAs will be considered as well. Different types of uncertainties will be included in the future work, like unexpected events (e.g. machine breakdowns, worker unavailability) and inaccurate information (e.g. stochastic processing times).